# Data Scoping: Effectively Learning the Evolution of Generic Transport PDEs

### v1

@@ 3.4 Data scoping
4 Numerical Experiments
In this section, we evaluate the proposed data scoping method in 2D Burgers’ equations for fluid dynamics and 3D heat transfer equations. We also report additional results for the 2D mass transport equation, which can be found in Appendix. C.1. By implementing the FNO and CNNs models with and without our data scoping method, its effects are revealed. The data generation processes are generalized here and detailed in Appendix C. The FNO and CNN models each have 4 layers with 20 hidden dimensions. The Adam optimizer [12] is employed with a learning rate of $0.0001$. Normalized $L_{2}$ is used as training and testing loss and the $R^{2}$ score is applied for validation metric. For the 2D problems, various window sizes were examined and their influences on the ML model performance were studied. For the 3D problem, it was determined that a window size of $7$ provides the best performance. The code is publicly available on  [^1] Google drive.
4.1 2D Burgers’ equation
The Burgers’ equation is also one of the most representative PDEs representing a convection-diffusion scheme. The solution of such an equation displays certain interesting physical phenomena including temporal wave propagation, shock wave formulation, and viscous-related energy dissipation. Approximating these complex dynamics is a challenging task for a machine learning model with no a priori information about the underlying physics, and therefore makes it a good testing case for validating model performance. Details of the physical equations and data generation methods are discussed in Appendix. C.2.
Examples of decomposed model prediction and reconstruction on Burgers’ equation are reported in Figure .6. The interesting finding is that while the solution of Burgers’ equation, compared with mass transport, contains much more complicated and localized physical dynamics, the data scoping method is still able to reconstruct smooth and continuous physical dynamics even though the dynamics inside each regional sub-domain greatly vary from each other.
Figure 6: Decomposed and reconstructed model prediction on solving the time sequential Burgers’ equation. All displayed physical properties are normalized and dimensionless.
4.2 Heat transfer in metal additive manufacturing
Thermal simulations for metal additive manufacturing (AM) are important in multiple stages of product development that involve AM processes, including part design, process planning, process monitoring, and process control [29, 23, 4, 5]. They are the typical examples where the traditional numerical methods are too time-consuming while the data-driven ML models are difficult to generalize to the situations not included in the training data. We test the capability of our data scoping method in improving the generalizability of the ML models for thermal simulations of the AM process. The type of generic transport PDEs in this problem is the heat transfer PDEs. The PDEs formulation and the data generated process are discussed in Appendix C.3 To study the geometric generalizability of the ML models, we collected the simulation data of 10 parts with various geometries. Figure 14 shows the 10 parts and snapshots of the thermal simulations. A 10-fold leave-one-outside cross-validation (LOOCV) is performed to evaluate the $R^{2}$ accuracy of the ML models over the geometry not included in the training data. At one round of LOOCV, the data from 9 parts are assembled as training and test data with a 9:1 ratio, and the data from the rest of one part is used for validation. Examples of the prediction results can be found in Figure 15.
5 Results and Discussion
To understand how the localized operator behaves under the data scoping, we perform a sequence of numerical experiments of different solution frequencies as stated in Section C.1. We examine the reconstructed $R^{2}$ accuracy of the solved solution field by FNO in correlation with the decomposed domain size and the character frequency of the solution field and report the result in Figure 7.
The frequency characterization of the domain reflects the speed of the information travels in the system. The higher the frequency, the faster the information travels and the larger the local-dependent region of the system becomes. For a fixed window size, as the frequency increase, the local-dependent region will gradually outgrows it and the model would be given less than required information to make the prediction at some point. That is why the accuracy drops dramatically as frequency increases for the smallest window size. It can also be explained in the view of frequency. A domain decomposition is a multiplication between the original solution function and a designated window function. Such multiplication therefore implements a frequency cutoff that adds high-frequency components to the Fourier domain of the original solution function. The higher the frequency of the original solution, the more divergence it will get when restricted to a localized area. To alleviate the effects of the frequency cut-off, we can increase the window size. However, as shown in Figure 7, a larger window size does not necessarily lead to better results. We see better solution quality as the decomposition window size increases from 6 to 10, however in high-frequency regions ($f>1$) as the window size gets larger than 12, the averaging accuracy obtained starts to decrease. It could be explained that the increased window size brings information not relevant to the local physical evolution, which is essentially noise so that it hinders the model from capturing the real physical pattern. It implies that there exists an optimal window size for a specific generic transport system. The results of the experiments indicate that the character frequency of the domain plays an important role in determining the window size.
The data scoping method can help accelerate the convergence of the ML models for generic transport problems by limiting the scope of input data. Figure 8 shows the average test loss over the 10-fold LOOCV of the CNNs and FNOs trained for the AM temperature prediction dataset with and without data scoping. The training processes with data scoping can reduce the test loss faster than the processes without data scoping. These results confirm our assumption that the coupling between the expressiveness and local dependency of deep learning architecture may expand the scope of input data beyond the travel range of the physics information, which can result in sluggish convergence. On the other hand, strictly limiting the scope of input data can speed up the training process.
The effects of the data scoping may differ depending on the type of ML models. As we can see in Figure 8, the FNOs gain a much larger improvement in the convergence rate than CNNs do. Such a difference might be originated from the local dependency of their linear operators. The linear operator of FNOs is the convolution in Fourier space, which involves the integral over the whole domain. So, the linear operator of FNOs is not local-dependent. Our data scoping method can make FNOs local-dependent while not influencing their powerful expressiveness in capturing the local physical patterns. On the other hand, the linear operator of CNNs is a kind of local convolution whose scope is prescribed by its kernel size. While the deep learning architecture can weaken its local dependency, CNNs are still more local-dependent than FNOs. Therefore, the improvements brought by the decoupling of the expressiveness and local dependency are more significant on FNOs than CNNs.
Figure 8: The average test loss with 95 % confidence intervals over the 10-fold LOOCV of the ML models trained for the AM temperature prediction dataset with and without data scoping. “DS” stands for data scoping. The stark difference between the effects of the data scoping method on CNN and FNO might be caused by the local dependency of their fundamental linear operator.
It should also be noted that the data scoping method can improve the generalizability of the ML models. Figure 9 shows the validation $R^{2}$ of the ML models over the 10-fold LOOCV of the AM temperature prediction dataset. The part index in the horizontal axis indicates the 10 parts with different geometries that are not included in the training and test of each validation round. So, the validation $R^{2}$ reveals the geometric generalizability of the models. As we can see, the data scoping method enhances the accuracy of all the validation data for both of the models. On average, CNNs are improved by 21.7 %, and FNOs are improved by 38.5%. This result supports our assumption that more data does not always beget better results. Especially, for the generic transport problems, the information outside of the local-dependent region is irrelevant to the local physical evolution. This extra information is essentially noises, which hinders the ML model from capturing the real physical patterns in the data. Therefore, limiting the scope of input data can effectively filter out the noises and help the ML models capture the real physical patterns contained in the data.
Figure 9: The data scoping method improves the geometric generalizability of the ML models for temperature prediction during AM processes. There are 10 rounds of leave-one-out cross-validation. In each round, the data of one geometric part is taken out and the ML models are trained on the data of the other 9 parts. Then, the ML models are validated over the data of the part not included in the training. “DS” stands for data scoping.
As analyzed in Section B.2, the time cost of the domain decomposition and window patching algorithms is linear with the largest block number in all dimensions $B_{max}$. A time cost experiment that collects the time cost of the two algorithms under different $B_{max}$ is conducted to demonstrate the linear time cost, as shown in Figure 10.
6 Conclusion and Future Direction
In this paper, we reveal the incompatibility between the deep learning architecture and the generic transport problems. We prove that the local-dependent region of deep learning models expands inevitably as the number of layers increases. On one hand, the expanded local-dependent region complicates the input data and introduces noise, which detrimentally impacts the convergence rate and generalizability of the models. On the other hand, the expressiveness of the ML models largely relies on the number of layers, so limiting the number of layers would weaken the performance of the models as well. Such a dilemma is caused by the coupled expressiveness and local-dependent property of deep learning architecture. To decouple the two, we propose an efficient data-scoping method. Through the numerical experiments over the data generated by three typical generic transport PDEs, we analyzed the properties of our data scoping method (e.g. the relationship among window size, system frequency, and error), and demonstrated its capabilities in accelerating convergence and enhancing generalizability of the ML models.
This approach has only been implemented on structured data now, but the idea has the potential for extension to unstructured data, like graphs. Since each partition is independent, parallel computation could accelerate the prediction integration significantly. By doing so, the scalability of our method, by computing each partition and each window independently, can be exploited and could have the potential for complex, large-scale generic transport problems.
Impact Statements
This paper presents work that seeks to broaden the application of machine learning methods in real-world engineering scenarios by improving the efficiency and robustness of neural network models for predicting physical properties. The authors do like to stress that, misusing the proposed method in safety-critical engineering simulations without sufficient validation can cause unexpected damage and negative social impact.
Appendix A Proofs
We prove Theorem 1 here.
Lemma 1.
Proof.
$d(z,x)\leq d(z,y_{0})+d(y_{0},x).$ (7)
Since $d(z,y_{0})<\delta_{2}$ and $d(y_{0},x)<\delta_{1}$, we have
$d(z,x)<\delta_{1}+\delta_{2}.$ (8)
So, $z\in U(x,\delta_{1}+\delta_{2})$.
$\xi=\sup d(z,y)+\sup d(y,x).$ (9)
Since $\sup d(y,x)=\delta_{1}$ and $\sup d(z,y)>\delta_{2}$, we have
$\xi>\delta_{1}+\delta_{2},$ (10)
Now we can prove Theorem 1.
Proof.
Since the non-linear activation function $\sigma$, and the linear project mapping $P$ and $Q$ do not influence the size of local-dependent region, we can simplify the expression of a neural operator as
$v_{i+1}=K_{\phi}(v_{i}),$ (11)
[…]
Appendix B Formulation of Data Scoping Method
B.1 Domain expansion
@@
B.2 Domain decomposition
Input data for the ML model is formed as batches. A batch of structured data can be represented as a tensor with size $(N_{b},N_{1},..,N_{d},N_{c})$ where $N_{b}$ is the batch number and $N_{c}$ is the channel dimension which is determined by the dimension of physical properties at a position. Given the window size $(W_{1},...,W_{d})$, our method aims to decompose the whole batch of data into a new tensor with size $(N_{b}*B_{i}*...*B_{d},W_{1},...,W_{d},N_{c})$. Appendix 1 shows the algorithm. Figure 4 illustrates a 2D example. In this example, we are given a batch of 2D data with size $(4,9,9,1)$ and the window size is $(3,3)$. The block number is $(3,3)$. After a sequence of splitting and stacking operations, the batch of the whole domain data is converted to a batch of the windows in the shape of $(36,3,3,1)$. For 2D data, there is twice splitting and twice stacking, while for 3D data, there is thrice splitting and thrice stacking. The time complexity of the splitting and stacking of the array data is $O(B_{max})$ where $B_{max}$ is the maximal among $B_{i},i=1,...,d$. The batch of the windows is then input into an ML model and the physical properties at the center of the windows are predicted as shown in Figure 4 (g). Then the window patching algorithm detailed in Appendix 2, a reverse of the decomposition operation, is followed to recover the batch of the whole domain from the batch of the windows. The window patching algorithm consists of the same number of splitting and stacking operations as the decomposition algorithm does, whose time complexity is also $O(B_{max})$. So the time complexity of the total algorithm is $O(B_{max})$.
B.3 Prediction integration
As shown in Figure 4 (h) the prediction over one decomposition can only give us the physical properties at the centers of the windows, to get the complete prediction over the whole domain, we need multiple decompositions which ensures that all the positions are the centers of some windows. Appendix 3 details the prediction integration algorithm. Figure 5 illustrates the prediction integration algorithm with a 2D example. Figure 5 (a) shows the 2D grid domain. The blank zone near the boundary indicates the padding zeros. The original domain size is $(9,9)$ and the expanded domain size is $(11,11)$. With $(3,3)$ window size, the block number for one decomposition is $(3,3)$. As each decomposition can be used to predict the center of all its windows, we need $3\times 3$ different decompositions of the domain which can cover all the positions as shown in Figure 5 (b) and (c). In general case, it is required to have $\prod_{i=1}^{d}W_{i}$ decompositions to cover the whole domain. The window size $W_{i}$ is a small number compared with the domain dimension. Since the predictions over the different decompositions are independent, they can be calculated in parallel. Therefore, the time complexity of the prediction integration algorithm is the constant multiple of the time complexity of the ML model inference.
@@ B.3 Prediction integration
Appendix C Data generation
C.1 Mass transport equation
The transport of mass can be seen as one of the most fundamental PDEs with variation in both time and space. It also enjoys the benefit of having a fully closed mathematical solution, and that the problem can be carefully constructed to show the solution field of different character frequencies. We test the data scoping method’s representation ability on solution functions of different frequencies to understand the decomposition method’s performance on questions including how small the decomposition can get, and how wide the frequency range the decomposition can capture before the model starts to lose accuracy due to domain cut-offs in Section. 5.
A typical mass transport equation can be expressed as Equation 15:
@@ C.1 Mass transport equation
for any initial condition $\mathbf{u}_{0}$, transport speed $c$ and given temporal stamp $t$. We can thus construct the frequency of our solution by determining the frequency of the initial condition. Variants of the solution frequency are shown in Figure. 11
@@
Examples of decomposed model prediction and reconstructions on mass transport equations using a solution frequency of 5 and window size of 10 are reported in Figure .12. We see that The model can achieve very high correspondence with physical dynamics even when trained on small windows that reduced the original data size by more than 36 times (6 times on each dimension). The reconstruction method applied in Figure. 5 also helped to smooth out the entire solution field, suggesting the potential of the model application in reconstructing physical dynamics at minimum cost.
Figure 12: Decomposed and reconstructed model prediction on solving the time sequential mass transport equation. All displayed physical properties are normalized and dimensionless.
C.2 Burgers’ equation
In our experiment, we implement the viscous version of Burgers’ equation as described by Eq. 17:
@@ C.2 Burgers’ equation
We solve Eq. 17 with $\nu=0.01Pa\cdot s$, and a time step of $0.1s$ for a total of 10 seconds with randomly initialized velocity Gaussian distribution as the initial condition. The simulations of the Burgers’ equation under different initial conditions are performed in FEniCS on a 2D mesh of $80\times 80$ elements per unit. Four Burgers’ equation solutions computed in four different initial velocity distributions are shown in Fig. 13.
@@
C.3 AM temperature data
Let $\Omega$ denote a domain, and $\partial\Omega$ its boundary. $\partial\Omega_{\mathrm{H}}$ represents the part at which heat is transferred to the surroundings with constant temperature $T_{\infty}$, and $\partial\Omega_{\mathrm{D}}$ the part at which the temperature is fixed at $T_{\mathrm{D}}$. The temperature evolution within $\Omega$ is governed by the heat transfer PDEs:

### v2

@@ 4.1 Method formulation
4.2 Domain decomposition
Input data for the ML model is formed as batches. A batch of structured data can be represented as a tensor with size $(N_{b},N_{1},..,N_{d},N_{c})$ where $N_{b}$ is the batch number and $N_{c}$ is the channel dimension which is determined by the dimension of physical properties at a position. Given the window size $(W_{1},...,W_{d})$, our method aims to decompose the whole batch of data into a new tensor with size $(N_{b}*B_{i}*...*B_{d},W_{1},...,W_{d},N_{c})$. Algorithm 1 shows the algorithm. Figure 4 illustrates a 2D example. In this example, we are given a batch of 2D data with size $(4,9,9,1)$ and the window size is $(3,3)$. The block number is $(3,3)$. After a sequence of splitting and stacking operations, the batch of the whole domain data is converted to a batch of the windows in the shape of $(36,3,3,1)$. For 2D data, there is twice splitting and twice stacking, while for 3D data, there is thrice splitting and thrice stacking. The time complexity of the splitting and stacking of the array data is $O(B_{max})$ where $B_{max}$ is the maximal among $B_{i},i=1,...,d$. The batch of the windows is then input into an ML model and the physical properties at the center of the windows are predicted as shown in Figure 4 (g). Then the window patching algorithm detailed in Algorithm 2, a reverse of the decomposition operation, is followed to recover the batch of the whole domain from the batch of the windows. The window patching algorithm consists of the same number of splitting and stacking operations as the decomposition algorithm does, whose time complexity is also $O(B_{max})$. So the time complexity of the total algorithm is $O(B_{max})$.
4.3 Domain expansion
@@
4.4 Prediction integration
As shown in Figure 4 (h) the prediction over one decomposition can only give us the physical properties at the centers of the windows, to get the complete prediction over the whole domain, we need multiple decompositions which ensures that all the positions are the centers of some windows. Algorithm 3 details the prediction integration algorithm. Figure 5 illustrates the prediction integration algorithm with a 2D example. Figure 5 (a) shows the 2D grid domain. The blank zone near the boundary indicates the padding zeros. The original domain size is $(9,9)$ and the expanded domain size is $(11,11)$. With $(3,3)$ window size, the block number for one decomposition is $(3,3)$. As each decomposition can be used to predict the center of all its windows, we need $3\times 3$ different decompositions of the domain which can cover all the positions as shown in Figure 5 (b) and (c). In general case, it is required to have $\prod_{i=1}^{d}W_{i}$ decompositions to cover the whole domain. The window size $W_{i}$ is a small number compared with the domain dimension. Since the predictions over the different decompositions are independent, they can be calculated in parallel. Therefore, the time complexity of the prediction integration algorithm is the constant multiple of the time complexity of the ML model inference.
@@ 4.4 Prediction integration
4.5 Window size determination
We define $L_{c}$ for hyperbolic PDEs in a similar way to the Courant–Friedrichs–Lewy (CFL) condition [38] in computational fluid dynamics which is a necessary condition for convergence while solving hyperbolic PDEs by guaranteeing the subdomain contains enough information about the flow of information. CFL condition specifies an upper bound for the time step $\Delta t$ with Courant number $C$ as
$C=\frac{c\delta t}{\delta x}\leq C_{max},$ (9)
where $c$ is the transport velocity, $\delta x$ is the element length of the discretization, $C_{max}$ is decided by experience which could range between 0.1 to 100 according to different solvers.
Similarly, we can define a lower bound for the distance the physical information travels with a fixed $\Delta t$, which is called the characteristic length $L_{c}$. $L_{c}$ can be determined according to the physical coefficients relevant to PDEs. For example, for mass transport PDEs, we have $L_{c}=c\Delta$ where $c$ is the mass transport speed whose base unit is $[L/T]$. For heat transfer PDEs, we have $L_{c}=\sqrt{\alpha\Delta}$ where $\alpha$ is the thermal diffusivity whose base unit is $[L^{2}/T]$. For Burger’s equation, there are two terms related to momentum transport, the convection and the diffusion terms, which have their own characteristic lengths respectively. For the convection term, we have $L_{c}^{conv}=u\Delta t$, where $u$ is the fluid velocity, which is similar to mass transport speed. For the diffusion term, we have $L_{c}^{diff}=\sqrt{\nu\Delta t}$, where $\nu$ is the diffusion coefficient, which has the same base unit as temperature diffusivity $\alpha$. We determine the characteristic length for Burger’s equation as $L_{c}=\max(L_{c}^{conv},L_{c}^{diff})$. The element length $\delta x$ of the discretization in our numerical simulation data is in the same scale of $L_{c}$.
From our experience, when the window size is in the scale of $\backsim 10L_{c}$, DDELD usually has the best effects in improving the model’s prediction accuracy. Given a specific dataset, we could not specify the optimal window size without a hyper-parameter tuning process because of two reasons. Firstly, the physical coefficients used to determine the characteristic lengths cannot precisely reflect the information traveling speed for all the points in the domain as the speed is affected by many other factors. For example, for mass transport, besides $c$, the transport speed is also affected by the gradients of mass concentration. Secondly, the initial and boundary conditions of the data would also affect the features of PDEs. Like all the data-driven methods, the hyper-parameters of the models could only be tuned through experience for data with various features.
What we find through experiments, however, indicates an alternate approach to finding a desired window size by analyzing the frequency distribution of the solution if we have some knowledge of it a priori. We find that data-driven methods for learning mappings between PDE operators are often more sensitive to the structure of the initial condition (or generally the input function to the model) than the coefficients in the PDE itself, see Appendix for a detailed explanation. This highlights the possibility of a frequency-based analysis for the choice of window size for PDEs. We give the following theorem as a guideline for the choice of window sizes for generic physical data with a clear frequency bias in Theorem 2 and appended the proof in Appendix A.
Theorem 2.
Define $u\in L_{1}(\mathbb{R})$ and let $\hat{u}$ be the Fourier transform of $u$ with a clear frequency bandwith of $\text{supp}(\hat{u})\subseteq[-B,B]$. If a set of observations of discretized values of $u$ per unit length $\{u(x_{i})|i=1,2,3,\cdots,N\}$ are available, and we obtain a local dependency region based on Equation 3, then the minimum number of points required for the local dependency region to retain frequency character would be given by: $L_{c}=\lceil{\frac{N+1}{2B}}\rceil$.
The experimental results regarding the influence of window size and information frequency over the model’s prediction accuracy is shown in Section 7.3.
5 Data Generation
In this section, we describe the generating process of the datasets used to evaluate the effects of DDELD, including 2D mass transport equation, 2D Burger’s equation, 2D isotropic turbulence in fluid dynamics, and 3D temperature transferring data.
5.1 Mass transport equation
The transport of mass can be seen as one of the most fundamental PDEs with variation in both time and space. It also enjoys the benefit of having a fully closed mathematical solution, and that the problem can be carefully constructed to show the solution field of different character frequencies. We test the DDELD’s representation ability on solution functions of different frequencies to understand the decomposition method’s performance on questions including how small the decomposition can get, and how wide the frequency range the decomposition can capture before the model starts to lose accuracy due to domain cut-offs in Section 7. A typical mass transport equation can be expressed as Equation 10:
@@ 5.1 Mass transport equation
for any initial condition $\mathbf{u}_{0}$, transport speed $c$ and given temporal stamp $t$. We can thus construct the frequency of our solution by determining the frequency of the initial condition. Variants of the solution frequency are shown in Figure 6.
@@
5.2 Burgers’ equation
The Burgers’ equation is also one of the most representative PDEs representing a convection-diffusion scheme. The solution of such an equation displays certain interesting physical phenomena including temporal wave propagation, shock wave formulation, and viscous-related energy dissipation. Approximating these complex dynamics is a challenging task for a machine learning model with no a priori information about the underlying physics, and therefore makes it a good testing case for validating model performance.
We implement the viscous version of Burgers’ equation as described by Equation 12:
@@ 5.2 Burgers’ equation
We solve Equation 12 with $\nu=0.01Pa\cdot s$, and a time step of $0.1s$ for a total of 10 seconds with randomly initialized velocity Gaussian distribution as the initial condition. The simulations of the Burgers’ equation under different initial conditions are performed in FEniCS on a 2D mesh of $80\times 80$ elements per unit. Four Burgers’ equation solutions computed in four different initial velocity distributions are shown in Figure 7.
@@
5.3 Isotropic turbulence
We performed testing and validation of the DDELD on simulation results with direct numerical simulation (DNS) on an isotropic turbulence scenario. The incompressible Navier-Stokes equation can be written as Equation 13:
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}=\nu\nabla^{2}\mathbf{u}-\frac{1}{\rho}\Delta p+\mathbf{f},~{}\mathbf{x}\in D$ (13)
where $\nu$ denotes the fluid viscosity, $\rho$ is the fluid density and $\mathbf{f}$ denotes the body force. We obtained DNS data from the Johns Hopkins Turbulence Database [39] (JHTDB) with a total of 5,028 sequential time steps solved with a pseudo-spectral solver. We present samples of the turbulent data in Figure 8
Figure 8: Solution of the isotropic turbulence at t=0s, t=0.02s, t=0.04s and t=0.06s respectively.
5.4 AM temperature numerical simulation
Thermal simulations for metal additive manufacturing (AM) are important in multiple stages of product development that involve AM processes, including part design, process planning, process monitoring, and process control [40, 41, 10, 3]. Because of the geometric complexity of the parts, AM thermal simulation is the typical scenario where the traditional numerical methods are too time-consuming while the data-driven ML models are difficult to generalize to the situations not included in the training data. So, it is critical to increase the geometric generalizability of data-driven models with limited data. Here, we generated an AM temperature dataset including the temperature histories of 10 parts in different geometries to evluate the effects of DDELD in improving the geometric generalizability of data-driven models. The 10 geometries are randomly generated by SkexGen [42], which contains various common mechanical features, such as holes, ribs, and pillars. Figure 9 shows the examples of the temperature data.
Figure 9: AM temperature prediction dataset. (a) The 10 parts with various geometries. (b) The temperature histories of the AM process that built the 10 parts are generated.
The heat transfer PDEs for AM process is formulated as follows. Let $\Omega$ denote a domain, and $\partial\Omega$ its boundary. $\partial\Omega_{\mathrm{H}}$ represents the part at which heat is transferred to the surroundings with constant temperature $T_{\infty}$, and $\partial\Omega_{\mathrm{D}}$ the part at which the temperature is fixed at $T_{\mathrm{D}}$. The temperature evolution within $\Omega$ is governed by the heat transfer PDEs: